Flexible Overhead Budget Leno Manufacturing Company prepared the following factory overhead cost budget for the Press Department for October of the current year, during which it expected to require 16,000 hours of productive capacity in the department: Variable overhead cost: Indirect factory labor $115,200 Power and light 6,880 Indirect materials 35,200 Total variable overhead cost $157,280 Fixed overhead cost: Supervisory salaries $55,050 Depreciation of plant and equipment 34,600 Insurance and property taxes 22,020 Total fixed overhead cost 111,670 Total factory overhead cost $268,950 Assuming that the estimated costs for November are the same as for October, prepare a flexible factory overhead cost budget for the Press Department for November for 14,000, 16,000, and 18,000 hours of production. Round your interim computations to the nearest cent, if required. Enter all amounts as positive numbers. Leno Manufacturing Company Factory Overhead Cost Budget-Press Department For the Month Ended November 30 Direct labor hours 14,000 16,000 18,000 Variable overhead cost: Indirect factory labor $ $ $ Power and light Indirect materials Total variable factory overhead $ $ $ Fixed factory overhead cost: Supervisory salaries $ $ $ Depreciation of plant and equipment Insurance and property taxes Total fixed factory overhead $ $ $ Total factory overhead cost
Answer #1
Factory Overhead Cost Budget
Direct labor hours           14,000           16,000           18,000
Variable overhead cost
Indirect factory labor        100,800        115,200        129,600
Power and Light             6,020             6,880             7,740
Indirect material           30,800           35,200           39,600
Total variable factory overhead        137,620        157,280        176,940
Fixed Factory overhead cost
Supervisory Salaries           55,050           55,050           55,050
Depreciation of Plant and Equipment           34,600           34,600           34,600
Insurance and Property taxes           22,020           22,020           22,020
Total fixed factory overhead        111,670        111,670        111,670
Total factory overhead cost        249,290        268,950        288,610
Note: Variable overhead varies with direct labor hours, but fixed overhead cost remains constant

D 14000 16000 18000 =115200*B2/16000 =115200*C2/16000 =115200*D2/16000 =6880*B2/16000 =6880*C2/16000 =6880*D2/16000 =35200*B2